# Peaceful Anarcho-Accelerationism: Decentralized Full Automation for a Society of Universal Care

### v1

@@ 5 Empirical Evidence: Commons That Work
Founded in 1956 in the Basque Country, Mondragon encompasses approximately 70,000 worker-owners across some 90 cooperatives, a university, 15 research centers, and a social welfare system. With over € 11 billion in annual revenue, it is the seventh-largest business group in Spain (Whyte and Whyte, 1991). Worker-owners vote on major decisions (one person, one vote), elect management, and the pay ratio between highest and lowest earners is capped at 6:1 (compared to 399:1 in large US corporations (Mishel and Kandra, 2021)). Table˜5 presents comparative data on cooperative versus conventional enterprise performance across multiple metrics, demonstrating that the cooperative model is not merely viable but superior on most dimensions.
@@
Table˜6 summarizes the cross-case evidence. These are not marginal experiments. They are large-scale, long-duration, empirically verified systems that embody the principles of anarcho-accelerationism. They demonstrate that a world built on care, cooperation, and commons is not a fantasy; it is a pattern that already exists and can be replicated.
@@
“This is utopian.” The case studies in Section˜5 are not utopian: they are operational systems with decades of track record. Linux is not naive; it runs the internet. Mondragon is not utopian; it employs 70,000 people. Calling a working system “utopian” is not an argument; it is a refusal to engage with evidence. As Ursula K. Le Guin wrote: “We live in capitalism. Its power seems inescapable. So did the divine right of kings” (Le Guin, 2014).
@@ 6 Roadmap: Milestones, Assumptions, and Limitations
A framework without a path of implementation is merely an aspiration. In this section we propose a phased roadmap toward the care-centered society, make explicit the assumptions on which it depends, identify the hard constraints it must respect, and acknowledge its limitations honestly. Table˜7 summarizes the four phases.
@@
Phase I is already underway. The technologies described in the Liberation Stack (Section˜4) are mature, the governance models are proven (Section˜5), and the cost curves for renewable energy and robotic hardware are declining exponentially (Figure˜3). The task of this phase is integration: assembling the existing components into complete, self-sustaining community stacks and demonstrating that they work together. Concretely, this means deploying community solar microgrids with open-source energy management, connecting them to fablabs with robotic manufacturing capability, establishing agroecological food production augmented by autonomous farming robots, linking everything through community-owned mesh networks, governing the whole through open-source democratic platforms, and beginning to experiment with UDR at the local level: residents of pilot communities request goods and services from the productive commons rather than purchasing them on markets.
@@
This roadmap rests on several explicit assumptions. First, we assume that the exponential trajectory of AI and robotic capability continues, as predicted by the universal approximation theorem and DRL convergence results (Section˜2), and as empirically observed over the past two decades (Table˜1). If progress stalls or decelerates significantly, Phases III and IV will be delayed but not invalidated: the Liberation Stack provides value at every stage of automation, not only at full automation. Second, we assume that sufficient renewable energy can be deployed to power full automation. Current projections from IRENA (IRENA, 2023) and the exponential decline in solar costs (Figure˜3) support this assumption, but it requires sustained investment in energy storage, grid infrastructure, and materials recycling. Third, we assume that open-source AI remains competitive with proprietary models. The current ecosystem (LLaMA, Mistral, Stable Diffusion, and their successors) suggests that this is viable, but sustained community investment in training infrastructure is essential. Fourth, we assume that communities can develop effective commons governance at scale. The empirical evidence (Section˜5) supports this for systems up to millions of participants, but planetary-scale federation remains an open institutional design problem.
@@ 7 Conclusion and Future Work
We have proposed the concept of Universal Desired Resources (UDR) as a post-monetary design principle that supersedes Universal Basic Income, grounded in the observation that money becomes obsolete when robotic commons can produce anything on demand. We have established the ethical permissibility of delegating all instrumental labor to machines by demonstrating the independence of phenomenal consciousness from computational intelligence (Garrido-Merchán and Lumbreras, 2022; Garrido-Merchán, 2024; Garrido-Merchán and Blanco, 2022), and we have proposed that moral policy for the commons can be studied and refined through deep reinforcement learning over humanist reward functions (Garrido-Merchán and Lumbreras-Sancho, 2023). We have demonstrated that anarcho-accelerationism builds genuine bridges to every major intellectual tradition (Section˜3), proposed the Liberation Stack as a concrete layered architecture for sociotechnical sovereignty (Section˜4), presented empirical evidence that commons-based systems already operate at planetary scale (Section˜5), and outlined a phased roadmap with explicit assumptions, constraints, and limitations (Section˜6).
@@
Several lines of future work follow directly from this framework. First, the UDR concept requires formal economic modeling: what are the equilibrium dynamics of a post-monetary commons economy? How does a UDR system handle heterogeneous preferences, luxury goods, and positional goods? What institutional mechanisms prevent the re-emergence of informal markets and power asymmetries? These questions demand the tools of mechanism design, institutional economics, and agent-based simulation. Second, the moral policy framework (Garrido-Merchán and Lumbreras-Sancho, 2023) must be extended from stylized models to realistic multi-agent environments: how do we define reward functions that faithfully encode humanist values without imposing cultural uniformity? How do communities democratically negotiate the parameters of their moral policy? How do we handle moral disagreement across federated communities? Third, the roadmap requires empirical validation through pilot deployments: the Liberation Stack must be implemented and tested in real communities, measuring not only technical performance but social outcomes (well-being, equity, ecological impact, community cohesion). Fourth, the convergence of additive manufacturing, synthetic biology, and recursive AI self-improvement toward universal fabrication capacity requires sustained interdisciplinary research at the intersection of materials science, molecular biology, and artificial intelligence. Fifth, the consciousness argument must be continually revisited as our understanding of phenomenal experience deepens: if computational architectures that give rise to consciousness are ever identified, the ethical framework must be updated accordingly.

### v2

@@ 6 Empirical Evidence: Commons That Work
Founded in 1956 in the Basque Country, Mondragon encompasses approximately 70,000 worker-owners across some 90 cooperatives, a university, 15 research centers, and a social welfare system. With over € 11 billion in annual revenue, it is the seventh-largest business group in Spain (Whyte and Whyte, 1991). Worker-owners vote on major decisions (one person, one vote), elect management, and the pay ratio between highest and lowest earners is capped at 6:1 (compared to 399:1 in large US corporations (Mishel and Kandra, 2021)). Table˜7 presents comparative data on cooperative versus conventional enterprise performance across multiple metrics, demonstrating that the cooperative model is not merely viable but superior on most dimensions.
@@
Table˜8 summarizes the cross-case evidence. These are not marginal experiments. They are large-scale, long-duration, empirically verified systems that embody the principles of anarcho-accelerationism. They demonstrate that a world built on care, cooperation, and commons is not a fantasy; it is a pattern that already exists and can be replicated.
@@
“This is utopian.” The case studies in Section˜6 are not utopian: they are operational systems with decades of track record. Linux is not naive; it runs the internet. Mondragon is not utopian; it employs 70,000 people. Calling a working system “utopian” is not an argument; it is a refusal to engage with evidence. As Ursula K. Le Guin wrote: “We live in capitalism. Its power seems inescapable. So did the divine right of kings” (Le Guin, 2014).
@@ 7 Roadmap: Milestones, Assumptions, and Limitations
A framework without a path of implementation is merely an aspiration. In this section we propose a phased roadmap toward the care-centered society, make explicit the assumptions on which it depends, identify the hard constraints it must respect, and acknowledge its limitations honestly. Table˜9 summarizes the four phases.
@@
Phase I is already underway. The technologies described in the Liberation Stack (Section˜5) are mature, the governance models are proven (Section˜6), and the cost curves for renewable energy and robotic hardware are declining exponentially (Figure˜3). The task of this phase is integration: assembling the existing components into complete, self-sustaining community stacks and demonstrating that they work together. Concretely, this means deploying community solar microgrids with open-source energy management, connecting them to fablabs with robotic manufacturing capability, establishing agroecological food production augmented by autonomous farming robots, linking everything through community-owned mesh networks, governing the whole through open-source democratic platforms, and beginning to experiment with UDR at the local level: residents of pilot communities request goods and services from the productive commons rather than purchasing them on markets.
@@
This roadmap rests on several explicit assumptions. First, we assume that the exponential trajectory of AI and robotic capability continues, as predicted by the universal approximation theorem and DRL convergence results (Section˜2), and as empirically observed over the past two decades (Table˜1). If progress stalls or decelerates significantly, Phases III and IV will be delayed but not invalidated: the Liberation Stack provides value at every stage of automation, not only at full automation. Second, we assume that sufficient renewable energy can be deployed to power full automation. Current projections from IRENA (IRENA, 2023) and the exponential decline in solar costs (Figure˜3) support this assumption, but it requires sustained investment in energy storage, grid infrastructure, and materials recycling. Third, we assume that open-source AI remains competitive with proprietary models. The current ecosystem (LLaMA, Mistral, Stable Diffusion, and their successors) suggests that this is viable, but sustained community investment in training infrastructure is essential. Fourth, we assume that communities can develop effective commons governance at scale. The empirical evidence (Section˜6) supports this for systems up to millions of participants, but planetary-scale federation remains an open institutional design problem.
@@ 8 Conclusion and Future Work
We have proposed the concept of Universal Desired Resources (UDR) as a post-monetary design principle that supersedes Universal Basic Income, and we have shown that UDR achieves the most radical intersectional intervention ever proposed: by eliminating the material basis of oppression, it simultaneously dissolves all axes of structural inequality identified by intersectional theory (Crenshaw, 1989; Collins, 2019), including class, race, gender, disability, sexuality, age, and nationality (Section˜4.1). Drawing on Maslow’s hierarchy of needs (Maslow, 1943), we have demonstrated that the Liberation Stack satisfies basic human needs universally, liberating every person to pursue self-actualization in what we term synthetic liberty: the positive freedom that emerges when commons-based infrastructure provides the material conditions for genuine autonomy, against the merely negative liberty of neoliberalism (Berlin, 1969; Sen, 1999) (Section˜4.2). We have proposed a progressive transition path from UBI to UDR that leverages capitalist technological development while building commons alternatives (Section˜4.3).
We have demonstrated that anarcho-accelerationism respects freedom of faith on principled epistemological and ontological grounds: the interior journey of the soul, as mapped by Saint Teresa of Avila and Saint Augustine (Lumbreras and Garrido-Merchán, 2024), lies categorically beyond the reach of any computational system, and full automation liberates time for the deepest forms of spiritual life (Section˜3). We have established the ethical permissibility of delegating all instrumental labor to machines by demonstrating the independence of phenomenal consciousness from computational intelligence (Garrido-Merchán and Lumbreras, 2022; Garrido-Merchán, 2024; Garrido-Merchán and Blanco, 2022), and we have proposed that moral policy for the commons can be studied and refined through deep reinforcement learning over humanist reward functions (Garrido-Merchán and Lumbreras-Sancho, 2023).
We have proposed the Liberation Stack as a concrete layered architecture for sociotechnical sovereignty, with its energy layer powered by frontier clean energy technologies (perovskite tandem solar, enhanced geothermal, nuclear fusion, iron-air batteries, green hydrogen) within an accelerationist ecological framework that achieves sustainability through technological abundance rather than degrowth (Section˜5). We have presented empirical evidence that commons-based systems already operate at planetary scale (Section˜6), and outlined a phased roadmap with explicit assumptions, constraints, and limitations (Section˜7).
@@
Several lines of future work follow directly from this framework. First, the UDR concept requires formal economic modeling: what are the equilibrium dynamics of a post-monetary commons economy? How does a UDR system handle heterogeneous preferences, luxury goods, and positional goods? What institutional mechanisms prevent the re-emergence of informal markets and power asymmetries? These questions demand the tools of mechanism design, institutional economics, and agent-based simulation. Second, the intersectional liberation thesis requires empirical validation: while the structural argument is sound, measuring the actual reduction of intersectional oppression in communities that adopt commons-based infrastructure is essential for refining the framework. Third, the moral policy framework (Garrido-Merchán and Lumbreras-Sancho, 2023) must be extended from stylized models to realistic multi-agent environments: how do we define reward functions that faithfully encode humanist values without imposing cultural uniformity? How do communities democratically negotiate the parameters of their moral policy? Fourth, the progressive transition from UBI to UDR requires pilot implementations in real communities, measuring not only technical performance but social outcomes (well-being, equity, ecological impact, community cohesion, intersectional equity, Maslow-level satisfaction). Fifth, the accelerationist ecological thesis must be validated against planetary boundaries: can frontier clean energy technologies, deployed at the scale required for full automation, operate within Earth’s biophysical limits? Sixth, the convergence of additive manufacturing, synthetic biology, and recursive AI self-improvement toward universal fabrication capacity requires sustained interdisciplinary research at the intersection of materials science, molecular biology, and artificial intelligence. Seventh, the consciousness argument must be continually revisited as our understanding of phenomenal experience deepens: if computational architectures that give rise to consciousness are ever identified, the ethical framework must be updated accordingly.